OpenStudy (anonymous):

g(x) = 3 sin(4x - pi) - 2 Find the y-intercept

jimthompson5910 (jim_thompson5910):

Plug in x = 0 to find the y-intercept g(x) = 3 sin(4x - pi) - 2 g(0) = 3 sin(4*0 - pi) - 2 g(0) = 3 sin(0 - pi) - 2 g(0) = 3 sin(-pi) - 2 g(0) = 3 (-sin(pi)) - 2 g(0) = 3(0) - 2 g(0) = -2 So we plugged in x = 0 as the input and we got g(0) = -2 as the output. Therefore, when x = 0, y = -2. The y-intercept as a point is (0,-2)

OpenStudy (anonymous):

Ya know... if I could just remember that f(x) is the same thing as y when I actually need it then maybe I wouldn't ask questions that make me feel dumb after I get an answer. Thanks for the help :)

jimthompson5910 (jim_thompson5910):

you're welcome and it's ok, just a silly mistake